OpenStudy (bohotness):

Simplify: 2 × 22 + 8 ÷ 2 - (5 + 1)

OpenStudy (anonymous):

\[\frac{ 2 × 22 + 8 }{ 2 - (5 + 1)}\]?

OpenStudy (anonymous):

or 2 × 22 + (8 ÷ 2 )- (5 + 1)

OpenStudy (anonymous):

44+4-6=?

OpenStudy (bohotness):

?

OpenStudy (compassionate):

Are you familiar with PEMDAS?

OpenStudy (bohotness):

42

OpenStudy (compassionate):

2 × 22 + 8 ÷ 2 - (5 + 1) 2 * 22 = 44 8/2 = 4 5 + 1 = 6 44 + 4 - 6 42 That's a simpler way to do it

OpenStudy (bohotness):

SO its 42